# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v5

@@ 4 Conclusion
Our interactive multiscale modeling framework, integrating FEA, MD, and DFT/AIMD, provides comprehensive insights into Li-CO2 battery electrochemistry. This approach accurately predicts cell performance while considering critical factors such as CO2 transport, cathode porosity, and Li2CO3/C deposition.
DFT/AIMD calculations reveal high electrical conductivity of the cathode catalyst, while MD simulations show favorable electrolyte properties. FEA results demonstrate an exponential decrease in discharge capacity with increasing current density, from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, primarily due to CO2 transport limitations and reduced cathode porosity. The Li2CO3/C deposit morphology transitions from larger structures at lower current densities to thin films at higher current densities, significantly impacting performance.
Our multiscale approach bridges atomic-level phenomena and cell-level performance, offering unique insights into spatiotemporal evolution of critical battery parameters and reducing the need for extensive prototyping. Future research should focus on enhancing CO2 transport, improving discharge product decomposition, and optimizing cathode designs with large surface areas, optimal pore sizes, and high porosity. This framework accelerates the development of high-performance Li-CO2 batteries and sustainable energy storage technologies by providing insights into phenomena difficult to observe experimentally in real-time.
@@ Appendix A. Supplementary material
Description of the experimental setup; a table of parameters used in the finite element analysis (FEA); a table of boundary conditions used in FEA; a table of initial values for independent and dependent variables used in FEA; plots showing the voltage-capacity curve during Li-CO2 battery discharge at various rates ranging from 0.2 mA/cm2 to 0.9 mA/cm2; plots showing the effect of initial values of relevant variables on the voltage-capacity curve during Li-CO2 battery discharge at a rate of 0.4 mA/cm2; plots illustrating changes in porosity and volume fraction of deposited species in the cathode throughout the discharge process at different current densities and discharge states; and plots showing the change in CO2 concentration within the cathode.

### v6

@@ 4 Conclusion
Our interactive multiscale modeling framework, integrating FEA, MD, and DFT/AIMD, provides comprehensive insights into Li-CO2 battery electrochemistry. This approach enables accurate prediction of cell performance by considering critical factors such as CO2 transport, cathode porosity, and Li2CO3/C deposition.
DFT/AIMD calculations reveal the high electrical conductivity of the cathode catalyst, while MD simulations demonstrate favorable electrolyte properties. FEA results show an exponential decrease in discharge capacity with increasing current density-from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2-primarily due to CO2 transport limitations and reduced cathode porosity. The morphology of Li2CO3/C deposits transitions from larger structures at lower current densities to thin films at higher current densities, significantly impacting performance.
Our multiscale approach bridges atomic-level phenomena and cell-level performance, offering unique insights into the spatiotemporal evolution of key battery parameters and reducing the need for extensive experimental prototyping. Future research should focus on enhancing CO2 transport, improving discharge product decomposition, and optimizing cathode design with large surface areas, optimal pore sizes, and high porosity. This framework accelerates the development of high-performance Li-CO2 batteries and sustainable energy storage technologies by providing insights into phenomena that are difficult to observe experimentally in real time. Moreover, the developed Li-CO2 battery continuum model at the cell level can be extended to study battery packs and thermal runaway. This modeling framework can also be adapted to other battery chemistries or used to simulate degradation, cycling behavior, and thermal effects, thereby broadening its applicability.
@@ Appendix A. Supplementary material
Background on catalyst/electrolyte developments for Li-CO2 battery; Description of the experimental setup; Computational details; Plots showing the voltage-capacity curve during Li-CO2 battery discharge at various rates ranging from 0.2 mA/cm2 to 0.9 mA/cm2; Plots showing the effect of initial values of relevant variables on the voltage-capacity curve during Li-CO2 battery discharge at a rate of 0.4 mA/cm2; Plots illustrating changes in porosity and volume fraction of deposited species in the cathode throughout the discharge process at different current densities and discharge states; and Plots showing the change in CO2 concentration within the cathode.